OpenStudy (hang254):

The two polygons shown below are similar. What is the length of side AD? 15 units 18 units 7 1/5 units 11 1/4 units

OpenStudy (campbell_st):

with out a diagram then is could be any of the 4

OpenStudy (hang254):

sorry, i forgot about that

OpenStudy (campbell_st):

using corresponding sides \[\frac{AD}{9} = \frac{10}{8}\] then\[AD = \frac{ 9 \times 10}{8}\] just solve

OpenStudy (hang254):

11.25, Thank you